OpenStudy (anonymous):

complete the following ordered pairs for the equation 6x + -5y = 20

OpenStudy (blurbendy):

which ordered pairs?

OpenStudy (anonymous):

(-5, ) , ( , -4) , ( , 2)

OpenStudy (blurbendy):

Okay, I'll do the first one 6(-5) + -5y = 20 -30 + -5y = 20 -5y = 50 y = -10 (-5, -10) Just substitute the value in the ordered pair for x or y in the equation Ordered pairs are as follows: (x, y)

OpenStudy (anonymous):

Thanks

OpenStudy (blurbendy):

np